OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = x^2 + 1, find f(g(3)).

OpenStudy (anonymous):

Sub g(x) into f(x) therefore: f(g(x)) = 2(x^2+1) + 3 = 2x^2 + 2 + 3 = 2x^2 + 5 then sub in 3 to answer

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

woah there. donwewege gave the right answer, but make sure you follow the guidance of the brackets. Helps you down the road in mathematics. When you understand algebra more fully (along with Calculus and whatever else you desire to learn), then you can discount brackets in many situations through your developed knowhow. answer should be as follows. f(x)= 2x + 3 g(x)= x^2+1 Find f(g(3): 2(3^2 + 1) +3 2(9+1) +3 2(10) +3 23

OpenStudy (anonymous):

thank you as well